# Beyond the Doors of Perception: Vision Transformers Represent Relations Between Objects

### v1

@@ Abstract
Though vision transformers (ViTs) have achieved state-of-the-art performance in a variety of settings, they exhibit surprising failures when performing tasks involving visual relations. This begs the question: how do ViTs attempt to perform tasks that require computing visual relations between objects? Prior efforts to interpret ViTs tend to focus on characterizing relevant low-level visual features. In contrast, we adopt methods from mechanistic interpretability to study the higher-level visual algorithms that ViTs use to perform abstract visual reasoning. We present a case study of a fundamental, yet surprisingly difficult, relational reasoning task: judging whether two visual entities are the same or different. We find that pretrained ViTs fine-tuned on this task often exhibit two qualitatively different stages of processing despite having no obvious inductive biases to do so: 1) a perceptual stage wherein local object features are extracted and stored in a disentangled representation, and 2) a relational stage wherein object representations are compared. In the second stage, we find evidence that ViTs can learn to represent somewhat abstract visual relations, a capability that has long been considered out of reach for artificial neural networks. Finally, we demonstrate that failure points at either stage can prevent a model from learning a generalizable solution to our fairly simple tasks. By understanding ViTs in terms of discrete processing stages, one can more precisely diagnose and rectify shortcomings of existing and future models.
@@ 1 Introduction
We analyze ViTs trained on two same-different tasks: an identity discrimination task, which constitutes the most basic instances of the abstract concept of same-different (and is most commonly studied in artificial systems) (Newport, 2021), and a relational match-to-sample task, which requires explicitly representing and manipulating an abstract concept of “same” or “different” (Cook & Wasserman, 2007; Geiger et al., 2023). See Figure 1 for examples of each. We relate the algorithms that models adopt to their downstream behavior, including compositional and OOD generalization [^1] Code is available at an anonymous github repo..
Our main contributions include:
@@
    We demonstrate that deficiencies in the perceptual stage can completely prevent models from learning abstract relational operations. By rectifying this first stage, these models can solve simple relational operations (i.e., discrimination) yet still struggle to learn more complex operations (i.e., RMTS).
@@ 2 Methods
Discrimination Task: We analyze a task involving the most basic instance of the same-different relation: the discrimination task. This task is well studied in machine learning (Kim et al., 2018; Tartaglini et al., 2023; Puebla & Bowers, 2022), and requires a mere comparison between two objects presented in a single image. It is also comparable to Identity-Match-to-Sample and Discrimination procedures in the cognitive sciences (Hochmann et al., 2017). Stimuli in the discrimination task consist of images containing two simple objects (See Figure 1, left). Each object may take on one of 16 different shapes and one of 16 different colors. The model is trained to distinguish whether these two tokens are the same along both color and shape dimensions or different along at least one of these dimensions. We train on 6,400 of these images. Notably, our stimuli are patch-aligned. Vision transformers tokenize images into square patches of N×$\times$×N pixels (in our case, $N=\{16,32\}$). We generate datasets such that objects reside completely within the bounds of a single patch (for the $N=32$ models) or within the bounds of exactly 4 patches (for the $N=16$ models). This simplification allows us to straightforwardly adopt techniques from mechanistic interpretability. To increase the difficulty of the task, tokens are randomly placed within patch boundaries, and Gaussian noise is sampled and added to the tokens.
@@
Relational Match-to-Sample Task: We also analyze a more abstract iteration of the same-different relation using a relational match-to-sample (RMTS) design. In this task, the model must generate an explicit representation of “sameness” and “difference”, and then operate over these representations (Martinho III & Kacelnik, 2016; Hochmann et al., 2017). Notably, though many species can solve the discrimination task, animals struggle to solve the RMTS task (Penn et al., 2008). Indeed, human children fail to solve RMTS tasks until around age 4 and only exhibit robust performance by approximately age 6 (Hochmann et al., 2017; Holyoak & Lu, 2021). Stimuli in the RMTS task contain 4 objects grouped in two pairs (See Figure 1, right). The “display” pair always occupies patches in the image’s top-left. The “sample” pair can occupy any other position. The task is defined as follows: for each pair, produce a same-different judgment (as in the discrimination task). Then, compare these judgments; if both pairs exhibit the same intermediate judgment (i.e., both pairs exhibit “same” or both pairs exhibit “different”), then the label is “same”. Otherwise, the label is different. Objects are identical to those found in the discrimination task, and they are patch-aligned in the same manner.
@@
Figure 2: Attention Pattern Analysis (a) CLIP Discrimination: Top: Heatmap showing the distribution of “local” (blue) vs. “global” (red) attention heads throughout a finetuned CLIP B/16 model. We see that local heads tend to cluster in the first several layers of the model, followed by a transition to global heads. Bottom: Line graph plotting the maximum proportion of attention paid by a single head to patches that are: Within the same object (within-object; WO), Within the other object that is being compared to the current patch’s object patch (within-pair; WP), in the background (BG). We see that WO values peak in early layers, followed by WP values, and finally, BG values. (b) From-Scratch Discrimination: We repeat the analysis in (a). This model does seem to contain local heads. (c) CLIP RMTS: We repeat the analysis for a CLIP model finetuned on RMTS. Top: Our results largely hold from (a). Bottom: We track another attention pattern: attention between pairs of objects (between pair; BP). We find that first, WO peaks, then WP, then BP, and finally BG. This accords with the hierarchical computations implied by the RMTS task. (d) DINO RMTS: We repeat the analysis in (c) for a DINO model and find no such pattern.
@@
Models: Tartaglini et al. (2023) demonstrated that CLIP-pretrained ViTs (Radford et al., 2021) can achieve high performance generalizing the same-different relation to out-of-distribution stimuli when fine-tuned on a same-different discrimination task. Thus, we primarily focus our analysis on this model [^2] Both 16-patch and 32-patch versions. Results for the 32-patch variant are presented in Appendix C.. In later sections, we compare to models pretrained using DINO (Caron et al., 2021) and ImageNet classification (Russakovsky et al., 2015; Dosovitskiy et al., 2020) objectives. We also train a randomly-initialized ViT model on each task. All models are finetuned on either a discrimination or RMTS task for 200 epochs, using the AdamW optimizer with a learning rate of 1E-6. We perform a sweep over learning rate schedulers (Exponential with a decay rate of $0.95$ and ReduceLROnPlateau with a patience of $40$). Models are selected by validation accuracy, and test accuracy is reported. Our results affirm and extend those presented in Tartaglini et al. (2023) – CLIP-pretrained ViTs perform extremely well on both tasks, achieving $\geq$97accuracy on a held-out test set. See Appendix B for results from all models.
@@ 3 Two-Stage Processing in ViTs
We now begin to characterize the internal mechanisms underlying the success of the CLIP ViT models. We discover that processing occurs in two distinct stages within the model: a perceptual stage, where the object tokens strongly attend to other tokens within the same object, and a relational stage, where tokens in one object attend to tokens in another object (or pair of objects).
@@
Methods – Attention Pattern Analysis: We explore the operations performed by the model’s attention heads, which “read” from particular patches and “write” that information to other patches (Elhage et al., 2021). In particular, we are interested in the flow of information within individual objects, between the two objects, and (in the case of RMTS) between two pairs of objects. We refer to attention heads that consistently exhibit within-object patterns across images as local attention heads and heads that attend to other tokens global attention heads, respectively. To classify an attention head as local or global, we score the head from 00 to $1$, where values closer to 00 indicate local operations and values closer to $1$ indicate global operations. To compute the score for an individual head, we collect attention patterns from the head on $500$ randomly selected “same” and “different” images ($1,000$ images total). Then, for each object in a given image, we compute the proportion of attention from the object’s patches to patches that do not comprise the object (excluding the CLS token) — including patches containing the other object(s) and non-object background tokens. [^3] We include attention from object to non-object tokens because we observe that models often move object information to a set of background register tokens (Darcet et al., 2023). See Appendix D. This yields two proportions, one for each object in the image. The attention head’s score for the image is the maximum of these two proportions. Finally, these scores are averaged across the images to produce the final score.
@@
Results: Attention head scores for CLIP fine-tuned on the discrimination and RMTS tasks are displayed in the heatmaps of Figure 2(a) and (c), respectively. The first half of these models is dominated by attention heads that most often perform local operations (blue cells). See Appendix D for examples of attention patterns. In the intermediate layers, attention heads begin to perform global operations reliably, and the deeper layers of the model are dominated by global heads. The prevalence of these two types of operations clearly demarcates two processing stages in CLIP ViTs: a perceptual stage where within-object processing occurs, followed by a relational stage where between-object processing occurs. These stages are explored in further depth in Sections 4 and  5 respectively [^4] We note that this attention pattern analysis is conceptually similar to Raghu et al. (2021), which demonstrated a general shift from local to global attention in ViTs. However, our analysis defines local vs. global heads in terms of objects rather than distances between patches..
@@
Figure 3: (a) Interchange interventions The base image exhibits the “different” relation, as the two objects differ in color. An interchange intervention extracts color information from a the intermediate representations generated by the same model run on a different image (the source image), and patches color information from the source image into the model’s intermediate representations of the base image. If successful, the intervened model will now return “same” when run on the base image. DAS is optimized to succeed at interchange interventions. (b) Disentanglement Results: We report the success of interchange interventions on shape and color across layers for CLIP B/16 finetuned on either the discrimination or RMTS task. We find that these properties are disentangled early in the model – one property can be manipulated without interfering with the other. The background is colored according to the heatmap in Figure 2(a).
@@
The line charts in Figure 2(a) and (b) show maximal scores of each attention head type (local and global) in each layer. Since (by construction) values closer to 00 indicate local heads, we plot these values subtracted from $1$. The global attention heads are further broken down into two subcategories for the discrimination task: object attention heads, whereby the tokens of one object attend to tokens associated with the object it is being compared to, and other attention heads whereby object tokens attend to background tokens. We add a third subcategory for RMTS: pair attention heads attend to tokens in the other pair of objects (e.g., a display object attending to a sample object). For both tasks, objects strongly attend to themselves during the first six layers, with a peak at layer 3. Throughout this perceptual stage, object heads steadily increase in prominence and peak in layer 6 (discrimination) or 5 (RMTS). In RMTS models, this object peak is followed by a pair peak, recapitulating the expected sequence of steps that one might use to solve RMTS. Notably, the object (and pair) peaks occur precisely where an abrupt transition from perceptual operations to relational operations occurs. Around layer 4, object attention to a set of background tokens begins to increase; after layer 6, object-to-background attention accounts for nearly all outgoing attention from object tokens. This suggests that processing may have moved into a set of register tokens (Darcet et al., 2023).
@@
Notably, this two-stage processing pipeline is not trivial to learn – neither a randomly initialized model trained on the discrimination task nor a DINO model trained on RMTS (Figure 2(b) and (d)) exhibit any obvious transition from local to global operations. We note that this two-stage processing pipeline loosely recapitulates the processing sequence found in biological vision systems: image representations are first formed during a feedforward sweep of the visual cortex, then feedback connections enable relational reasoning over these representations Kreiman & Serre (2020).
@@ 4 The Perceptual Stage
Attention between tokens is largely restricted to other tokens within the same object in the perceptual stage, but to what end? We find that these layers produce disentangled local object representations which encode shape and color. These properties are represented in separate linear subspaces within the intermediate representations of CLIP-pretrained ViTs.
Methods – Distributed Alignment Search: Distributed Alignment Search (DAS) is used to identify whether particular variables are causally implicated in a model’s computation (Geiger et al., 2024; Wu et al., 2024c) [^5] DAS has recently come under scrutiny for being too expressive (and thus not faithful to the model’s internal algorithms) when misused (Makelov et al. (2023), cf. Wu et al. (2024b)) Following best practices, we deploy DAS on the residual stream. We use the pyvene library for all DAS experiments (Wu et al., 2024a).. Given a neural network $M$, hypothesized high-level causal model $C$, and high-level variable v, DAS attempts to isolate a linear subspace $s$ of the residual stream states generated by $M$ that represents $v$ (i.e. $s$ takes on a value $s_{1}$ to represent $v_{1}$, $s_{2}$ to represent $v_{2}$, and so on). The success of DAS is measured by the success of counterfactual interventions. If $C(v_{1})=y_{1}$ and $C(v_{2})=y_{2}$, and $M(x)=y_{1}$ for some input $x$, does replacing $s_{1}$ with $s_{2}$ change the model’s decision to $y_{2}$?
Figure 4: (a) Novel Representations Analysis: Using trained DAS interventions, we can inject any vector into the shape or color subspaces that a model produces, allowing us to test whether the same-different operation can be computed over arbitrary vectors. We intervene on a “different”’ image – which differs only in its color property – by patching a novel color (an interpolation of red and black) into both objects in order to flip the decision to “same”. (b) Results: Results from performing the novel representations analysis using four methods for generating novel representations: adding observed representations, interpolating observed representations, per-dimension sampling using a distribution derived from observed representations, and finally sampling randomly from a mean 0, standard deviation 1 normal distribution. We see that the model’s same-different operation seems to generalize to vectors generated by adding (and to generalize somewhat to interpolated vectors) in early layers but not to sampling. The background is colored according to the heatmap in Figure 2(a).
@@
Results: We identify independent linear subspaces for color and shape in the intermediate representations produced in the early layers of a CLIP-pretrained ViT (See Figure3(b)). In other words, we can extract either color or shape information from one object and inject into another object. This holds for both discrimination and RMTS tasks. One can conclude that at least one function of the perceptual stage is to form disentangled local object representations, which are then used to solve same-different tasks. Notably, we see that local object representations are formed in the first few layers, becoming increasingly irrelevant deeper into the model. Indeed, intervening on the intermediate representations of object tokens at layers 5 and beyond results in chance intervention performance or worse.

### v3

@@ Abstract
Though vision transformers (ViTs) have achieved state-of-the-art performance in a variety of settings, they exhibit surprising failures when performing tasks involving visual relations. This begs the question: how do ViTs attempt to perform tasks that require computing visual relations between objects? Prior efforts to interpret ViTs tend to focus on characterizing relevant low-level visual features. In contrast, we adopt methods from mechanistic interpretability to study the higher-level visual algorithms that ViTs use to perform abstract visual reasoning. We present a case study of a fundamental, yet surprisingly difficult, relational reasoning task: judging whether two visual entities are the same or different. We find that pretrained ViTs fine-tuned on this task often exhibit two qualitatively different stages of processing despite having no obvious inductive biases to do so: 1) a perceptual stage wherein local object features are extracted and stored in a disentangled representation, and 2) a relational stage wherein object representations are compared. In the second stage, we find evidence that ViTs can sometimes learn to represent abstract visual relations, a capability that has long been considered out of reach for artificial neural networks. Finally, we demonstrate that failures at either stage can prevent a model from learning a generalizable solution to our fairly simple tasks. By understanding ViTs in terms of discrete processing stages, one can more precisely diagnose and rectify shortcomings of existing and future models.
@@ 1 Introduction
We analyze ViTs trained on two same-different tasks: an identity discrimination task, which constitutes the most basic instances of the abstract concept of same-different (and is most commonly studied in artificial systems) (Newport, 2021), and a relational match-to-sample task, which requires explicitly representing and manipulating an abstract concept of “same” or “different” (Cook & Wasserman, 2007; Geiger et al., 2023). See Figure 1 for examples of each. We relate the algorithms that models adopt to their downstream behavior, including compositional and OOD generalization [^1] Code is available here..
Our main contributions are the following:
@@
    We demonstrate that deficiencies in either the perceptual or relational stage can completely prevent models from learning abstract relational operations. By rectifying either stage, models can solve simple relational operations. However, both stages must be intact in order to learn more complex operations.
@@ 2 Methods
Discrimination Task. The discrimination task tests the most basic instance of the same-different relation. This task is well studied in machine learning (Kim et al., 2018; Puebla & Bowers, 2022; Tartaglini et al., 2023) and simply requires a single comparison between two objects. Stimuli in our discrimination task consist of images containing two simple objects (see Figure 1a). Each object may take on one of 16 different shapes and one of 16 different colors (see Appendix A for dataset details). Models are trained to classify whether these two objects are the “same” along both color and shape dimensions or “different” along at least one dimension. Crucially, our stimuli are patch-aligned. ViTs tokenize images into patches of $N\times N$ pixels ($N=\{14,16,32\}$). We generate datasets such that individual objects reside completely within the bounds of a single patch (for $N=32$ models) or within exactly 4 patches (for $N=16$ and $N=14$ models). Patch alignment allows us to adopt techniques from NLP mechanistic interpretability, which typically assume meaningful discrete inputs (e.g. words). To increase the difficulty of the task, objects are randomly placed within patch boundaries, and Gaussian noise is sampled and added to the tokens (see Appendix A). Models are trained on $6,400$ images.
We evaluate these models on out-of-distribution synthetic stimuli to ensure that the learned relations generalize. Finally, we evaluate a rather extreme form of generalization by generating a “realistic” dataset of discrimination examples using Blender, which include various visual attributes such as lighting conditions, backgrounds, and depth of field (see Appendix A.3).
@@
Relational Match-to-Sample (RMTS) Task. Due to the simplicity of the discrimination task, we also analyze a more abstract (and thus more difficult) iteration of the same-different relation using a relational match-to-sample (RMTS) design. In this task, the model must generate explicit representations of “sameness” and “difference”, and then operate over these representations (Martinho III & Kacelnik, 2016; Hochmann et al., 2017). Although many species can solve the discrimination task, animals (Penn et al., 2008) and children younger than 6 (Hochmann et al., 2017; Holyoak & Lu, 2021) struggle to solve the RMTS task. Stimuli in the RMTS task contain 4 objects grouped in two pairs (Figure 1b). The “display” pair always occupies patches in the top left of the image. The “sample” pair can occupy any other position. The task is defined as follows: for each pair, produce a same-different judgment (as in the discrimination task). Then, compare these judgments—if both pairs exhibit the same intermediate judgment (i.e., both pairs exhibit “same” or both pairs exhibit “different”), then the label is “same”. Otherwise, the label is “different.” Objects are identical to those in the discrimination task, and they are similarly patch-aligned.
@@
Figure 2: Attention Pattern Analysis. (a) CLIP Discrimination: The heatmap (top) shows the distribution of “local” (blue) vs. “global” (red) attention heads throughout a CLIP ViT-B/16 model fine-tuned on discrimination (Figure 1a). The $x$-axis is the model layer, while the $y$-axis is the head index. Local heads tend to cluster in early layers and transition to global heads around layer 6. For each layer, the line graph (bottom) plots the maximum proportion of attention across all $12$ heads from object patches to image patches that are 1) within the same object (within-object$=$WO), 2) within the other object (within-pair$=$WP), or 3) in the background (BG). The stars mark the peak of each. WO attention peaks in early layers, followed by WP, and finally BG. (b) From Scratch Discrimination: We repeat the analysis in (a). The model contains nearly zero local heads. (c) CLIP RMTS: We repeat the analysis for a CLIP model fine-tuned on RMTS (Figure 1b). Top: Our results largely hold from (a). Bottom: We track a fourth attention pattern—attention between pairs of objects (between pair$=$BP). We find that WO peaks first, then WP, then BP, and finally BG. This accords with the hierarchical computations implied by the RMTS task. (d) DINO RMTS: We repeat the analysis in (c) for a DINO model and find no such hierarchical pattern.
@@
Models. Tartaglini et al. (2023) demonstrated that CLIP-pretrained ViTs (Radford et al., 2021) can achieve high performance in generalizing the same-different relation to out-of-distribution stimuli when fine-tuned on a same-different discrimination task. Thus, we primarily focus our analysis on this model [^2] Both B/16 and B/32 versions. Results for the B/32 variant are presented in Appendix C.. In later sections, we compare CLIP to additional ViT models pretrained using DINO (Caron et al., 2021), DINOv2 (Oquab et al., ), masked auto encoding (MAE; He et al. (2022)) and ImageNet classification (Russakovsky et al., 2015; Dosovitskiy et al., 2020) objectives. We also train a randomly-initialized ViT model on each task (From Scratch). All models are fine-tuned on either a discrimination or RMTS task for 200 epochs using the AdamW optimizer with a learning rate of 1e-6. We perform a sweep over learning rate schedulers (Exponential with a decay rate of $0.95$ and ReduceLROnPlateau with a patience of $40$). Models are selected by validation accuracy, and test accuracy is reported. Our results affirm and extend those presented in Tartaglini et al. (2023)—CLIP and DINOv2-pretrained ViTs perform extremely well on both tasks, achieving $\geq$97accuracy on a held-out test set. Appendix B presents results for all models.
@@ 3 Two-Stage Processing in ViTs
We now begin to characterize the internal mechanisms underlying the success of the CLIP and DINOv2 ViT models. In the following section, we cover how processing occurs in two distinct stages within the model: a perceptual stage, where the object tokens strongly attend to other tokens within the same object, and a relational stage, where tokens in one object attend to tokens in another object (or pair of objects).
@@
Methods — Attention Pattern Analysis. We explore the operations performed by the model’s attention heads, which “read” from particular patches and “write” that information to other patches (Elhage et al., 2021). In particular, we are interested in the flow of information within individual objects, between the two objects, and (in the case of RMTS) between two pairs of objects. We refer to attention heads that consistently exhibit within-object patterns across images as local attention heads and heads that attend to other tokens global attention heads. To classify an attention head as local or global, we score the head from 00 to $1$, where values closer to 00 indicate local operations and values closer to $1$ indicate global operations. To compute the score for an individual head, we collect its attention patterns on $500$ randomly selected “same” and “different” images ($1,000$ images total). Then, for each object in a given image, we compute the proportion of attention from the object’s patches to any other patches that do not belong to the same object (excluding the CLS token)—this includes patches containing the other object(s) and non-object background tokens. [^3] We include attention from object to non-object tokens because we observe that models often move object information to a set of background register tokens (Darcet et al., 2023). See Appendix F. This procedure yields two proportions, one for each object in the image. The attention head’s score for the image is the maximum of these two proportions. Finally, these scores are averaged across the images to produce the final score.
@@
Results. Attention head scores for CLIP ViT-B/16 fine-tuned on the discrimination and RMTS tasks are displayed in the heatmaps of Figure 2 and  2c, respectively. The first half of these models is dominated by attention heads that most often perform local operations (blue cells). See Appendix F for examples of attention patterns. In the intermediate layers, attention heads begin to perform global operations reliably, and the deeper layers of the model are dominated by global heads. The prevalence of these two types of operations clearly demarcates two processing stages in CLIP ViTs: a perceptual stage where within-object processing occurs, followed by a relational stage where between-object processing occurs. These stages are explored in further depth in Sections 4 and  5 respectively. [^4] We note that this attention pattern analysis is conceptually similar to Raghu et al. (2021), which demonstrated a general shift from local to global attention in ViTs. However, our analysis defines local vs. global heads in terms of objects rather than distances between patches. We also find similar two-stage processing when evaluating on a discrimination dataset that employs realistic stimuli, suggesting that the patterns observed on our synthetic stimuli are robust and transferable (see Appendix A.3).
@@
Figure 3: (a) Interchange interventions: The base image exhibits the “different” relation, as the two objects differ in either shape (top) or color (bottom). An interchange intervention extracts {shape, color}information from the intermediate representations generated by the same model run on a different image (source), then patches this information from the source image into the model’s intermediate representations of the base image. If successful, the intervened model will now return “same” when run on the base image. DAS is optimized to succeed at interchange interventions. (b) Disentanglement Results: We report the success of interchange interventions on shape and color across layers for CLIP ViT-B/16 fine-tuned on either the discrimination or RMTS task. We find that these properties are disentangled early in the model—one property can be manipulated without interfering with the other. The background is colored according to the heatmap in Figure 2a, where blue denotes local heads and red denotes global heads.
@@
The line charts in Figure 2 show maximal scores of each attention head type (local and global) in each layer. Since values closer to 00 indicate local (i.e., within-object; WO in Figure 2) heads by construction, we plot these values subtracted from $1$. The global attention heads are further broken down into two subcategories for the discrimination task: within-pair attention heads, whereby the tokens of one object attend to tokens associated with the object it is being compared to (WP in Figure 2), and background attention heads, whereby object tokens attend to background tokens (BG in Figure 2). We add a third subcategory for RMTS: between-pair attention heads, which attend to tokens in the other pair of objects (e.g., a display object attending to a sample object; BP in Figure 2). For both tasks, objects strongly attend to themselves throughout the first six layers, with a peak at layer 3. Throughout this perceptual stage, within-pair heads steadily increase in prominence and peak in layer 6 (discrimination) or 5 (RMTS). In RMTS models, this within-pair peak is followed by a between-pair peak, recapitulating the expected sequence of steps that one might use to solve RMTS. Notably, the within-pair (and between-pair) peaks occur precisely where an abrupt transition from perceptual operations to relational operations occurs. Around layer 4, object attention to a set of background tokens begins to increase; after layer 6, object-to-background attention accounts for nearly all outgoing attention from object tokens. This suggests that processing may have moved into a set of register tokens (Darcet et al., 2023).
@@
Notably, this two-stage processing pipeline is not trivial to learn—several models, including a randomly initialized model trained on the discrimination task and a DINO model trained on RMTS (Figure 2b and d) fail to exhibit any obvious transition from local to global operations (See Appendix E for results from other models). However, we do find this pipeline in DINOv2 and ImageNet pretrained models (See Appendix D and E). We note that this two-stage processing pipeline loosely recapitulates the processing sequence found in biological vision systems: image representations are first formed during a feedforward sweep of the visual cortex, then feedback connections enable relational reasoning over these representations Kreiman & Serre (2020).
@@ 4 The Perceptual Stage
Attention between tokens is largely restricted to other tokens within the same object in the perceptual stage, but to what end? In the following section, we demonstrate that these layers produce disentangled local object representations which encode shape and color. These properties are represented in separate linear subspaces within the intermediate representations of CLIP and DINOv2-pretrained ViTs.
Methods — DAS. Distributed Alignment Search (DAS) (Geiger et al., 2024; Wu et al., 2024c) is used to identify whether particular variables are causally implicated in a model’s computation [^5] DAS has recently come under scrutiny for being too expressive (and thus not faithful to the model’s internal algorithms) when misused (Makelov et al. (2023), cf. Wu et al. (2024b)). Following best practices, we deploy DAS on the residual stream. We use the pyvene library (Wu et al., 2024a) for all DAS experiments.. Given a neural network $M$, hypothesized high-level causal model $C$, and high-level variable v, DAS attempts to isolate a linear subspace $s$ of the residual stream states generated by $M$ that represents $v$ (i.e. $s$ takes on a value $s_{1}$ to represent $v_{1}$, $s_{2}$ to represent $v_{2}$, and so on). The success of DAS is measured by the success of counterfactual interventions. If $C(v_{1})=y_{1}$ and $C(v_{2})=y_{2}$, and $M(x)=y_{1}$ for some input $x$, does replacing $s_{1}$ with $s_{2}$ change the model’s decision to $y_{2}$?
Figure 4: (a) Novel Representations Analysis: Using trained DAS interventions, we can inject any vector into a model’s shape or color subspaces, allowing us to test whether the same-different operation can be computed over arbitrary vectors. We intervene on a “different” image—differing only in its color property—by patching a novel color (an interpolation of red and black) into both objects in order to flip the decision to “same”. (b) Discrimination Results: We perform novel representations analysis using four methods for generating novel representations: 1) adding observed representations, 2) interpolating observed representations, 3) per-dimension sampling using a distribution derived from observed representations, and 4) sampling randomly from a normal distribution $\mathcal{N}(0,1)$. The model’s same-different operation generalizes well to vectors generated by adding (and generalizes somewhat to interpolated vectors) in early layers but not to sampled or random vectors. The background is colored according to the heatmap in Figure 2a (blue$=$local heads; red$=$global heads).
@@
Results. We identify independent linear subspaces for color and shape in the intermediate representations produced in the early layers of CLIP-pretrained ViTs (Figure 3b). In other words, we can extract either color or shape information from one object and inject it into another object. This holds for both discrimination and RMTS tasks. One can conclude that at least one function of the perceptual stage is to form disentangled local object representations, which are then used to solve same-different tasks. Notably, these local object representations are formed in the first few layers and become increasingly irrelevant in deeper layers; intervening on the intermediate representations of object tokens at layers 5 and beyond results in chance intervention performance or worse. DINOv2-pretrained ViTs provide similar results (Appendix D), whereas other models exhibit these patterns less strongly (Appendix I). We present a control experiment in Appendix H, which further confirms our interpretation of these results.